THE IONIZATION TOWARD THE HIGH-MASS STAR-FORMING REGION NGC 6334 I

\underline{J. MORALES}, University of Puerto Rico, Rio Piedras Campus, Physics Department, San Juan, Puerto Rico 00931; C. CECCARELI, Institut de Planétologie et d’Astrophysique de Grenoble (IPAG) UMR 5274, Grenoble, F-38041, France; L. OLMI, Osservatorio Astrofisico di Arcetri - INAF, Largo E. Fermi 5, I-50125, Firenze, Italy; D. LIS, California Institute of Technology, Pasadena, CA 91125, USA; R. PLUME, Department of Physics and Astronomy, University of Calgary, Calgary, AB T2N 1N4, Canada; P. SCHILKE, I. Physikalisches Institut der Universität zu Köln, Zülpicher Str. 77, 50937 Köln, Germany.

In this work we present high-resolution spectral line observations carried out with the HIFI (Heterodyne Instrument for the Far Infrared) instrument on board the Herschel Space Observatory (HSO) toward the so-called hot core NGC 6334 I, a region of high-mass star formation. From the Herschel/HIFI observations of the molecular tracers C^{18}O, C^{17}O, HCO^{+}, H^{13}CO^{+}, and N_{2}H^{+}, we determined the line parameters of each of the rotational transitions, with J_{up} ≥ 5. With these results, and using a non-Local Thermodynamic Equilibrium (LTE) Large Velocity Gradient (LVG) radiative transfer code, we modeled the spectral line emission of the molecular transitions in order to estimate the temperature, H_{2} density, source size, and CO, HCO^{+} and N_{2}H^{+} column density toward NGC 6334 I. We also derived the H_{2} column density as well as the relative abundances between the various molecular species by using the results from the LVG analysis. Finally, by modeling the chemical evolution of the source with a gas-phase reaction network, we obtained an estimate of the cosmic ray ionization rate toward the region. Based on our results, we conclude that there is an expansion of the envelope surrounding the hot core of NGC 6334 I, and the physical and kinematical properties of this expanding envelope are described in our analysis.